Please show your work: list out the N, I/Y, PV, PMT, and FV with correct signs. 2. Moon has been investing $2,500 quarterly for the past 10 years in an equity mutual fund. How much is the fund worth now assuming she has earned 8.5% compounded quarterly on her investment? I/Y PV PMT FV
Answer #1

Here we have considered cash outflow as negative and inflow as positive

N=10*4=40
I/Y=8.5/4=2.125
PV=-0
PMT=-2500

CPT FV = $155,165.18